OpenStudy (anonymous):

Solve for y: 10(1-2y) = -5(2y-1)

OpenStudy (austinl):

Distribute, terms with variables on one side, constants on the other. Divide.

OpenStudy (anonymous):

10(1-2y) = -5(2y-1) -2(1-2y) = (2y-1) 4y-2 = 2y-1 2y = 1 y = 1/2